# End-to-end Learned Visual Odometry with Events and Frames

### v2

@@ IV Experiments
Training. We train RAMP-VO on the synthetic TartanAir [5] dataset, with same train and test-split in [1], which we augment with events using VID2E [48]. During training, we follow the DPVO [1] training scheme to select and filter frames for training. Additionally, we also enforce a minimum of $1.2$M events between every pair of frames and remove additional frames if this condition is not satisfied. Finally, we feed RAMP-VO with both events and frames by interleaving two event stacks for every pair of selected frames. We create the first event stack by stacking the $600,000$ events received just before the mid-timestamp between the pair of images, and the second event stack by aggregating the $600,000$ events preceding the second frame. Since ground truth poses are only given for frames, we do not compute a loss for the mid-frame events. At inference time, we feed the model with asynchronous images and events, by creating a new event stack each time $M$ events are received. As the number of triggered events depends on the scene dynamics, with more events triggered with faster motion, the number of event stacks collected relative to a scene varies adaptively. For comparisons, we also test our model with the same strategy used during training, which we call synchronized since events are collected at regular rates.
@@
Datasets.  We use the Stereo DAVIS [49], EDS [2] dataset, and the TartanAir [5] test-split used in the ECCV 2020 SLAM competition to compare our method against the state-of-the-art. We use the TartanAir test-split used in [1] to ablate the crucial components of RAMP-VO architecture. Additionally, due to the limited availability of datasets with challenging lighting conditions, we introduce two new benchmarks that replicate lunar surface landings: Apollo landing and Malapert landing. These benchmarks feature rapid motion, high dynamic range, and textureless terrains.
Malapert landing.  Malapert landing consists of $20$ minutes of simulated data, divided into $2$ sequences of the Malapert south Moon region. We use the planets and satellites simulator PANGU  [^3] https://pangu.software/ to generate realistic descent trajectories, each $250$ km long in altitude and $40$ km in translation, featuring partial or complete darkness. Ground truth poses of the spacecraft’s center of mass are provided at $5$Hz, together with $640\times 480$ synchronized RGB images. We generate synthetic events using Vid2E [48] with default settings.
Apollo landing.  Apollo landing is a real dataset consisting of a total of $5$ minutes of recording, split into $6$ trajectories, featuring both vertical and lateral descent trajectories on a $260\times 260$ cm scale replica of the Apollo 17 landing site. Frames and events are recorded with a beam-splitter setup featuring a $20$Hz FLIR BFS-U3-12263c RGB camera and a Prophesee Gen4 event camera, similar to the one used in [2]. Poses are recorded with an OptiTrack motion capture system. We downsample frames and events to a resolution of $640\times 480$ before processing.
@@ IV-A Effects of RAMP blocks
Results: From Figure 4 (a), DPVO+events has the lowest performance among the methods we tested (AUC of $0.56$), highlighting that a trivial event-image fusion is not sufficient to exploit events. Moreover, our RAMP-VO encoder in synchronous mode outperforms the baseline that uses RAM-Net both when we use multiple scales, as in RAM-Net, but also when just one scale is used. The single-scale RAMP-VO (AUC of $0.71$) achieves a 11% increase over the RAM-Net encoder (AUC $0.64$), while the multi-scale version (AUC $0.81$) improves by 33% over RAM-Net. When events are processed asynchronously, using $M=250000$, the performance increases even more, reaching a $0.85$ AUC.
@@
Low Framerate VO: In this section, we demonstrate that our method is not only able of asynchronous processing, but it can also cope with frames provided at very low rates. Since our RAMP encoders build sensor agnostic features, we can exploit the faster data stream, like events voxels, to keep the feature embedding updated, and generate a consistent pose.
To test this capability, we design an experiment on the carwelding hard p003 and carwelding easy p007 sequences of the TartanAir, where we artificially reduce the framerate of the images, subsampling them, but keeping event stacks fixed at 20 Hz. We then evaluate the trajectory error against ground truth poses at 20 Hz for DPVO+events, RAMP-VO, and RAMP-VO with RAM-Net-like encoder, and report the results in Fig. 4 (c). Both the RAM-Net-like and RAMP encoders exhibit lower absolute trajectory errors compared to DPVO+events, particularly at lower framerates. This improvement stems from their ability to process asynchronous events in between frames. Notably, however, our RAMP encoder consistently outperforms the RAM-Net-like encoder, indicating its superior performance.
@@ IV-B Results on Space Data
DPVO [1] I 73,1 48,2 0,9 0,3 1,3
DPVO [1] I+E 40,1 34,0 0,9 0,2 1,1
RAMP-VO SS (Ours) I+E 1,1 9,4 0,7 0,2 1,0
RAMP-VO (Ours) I+E 0,6 4,3 0,8 0,2 0,9
TABLE II: Average Absolute Trajectory Error (cm) on Stereo Davis
Input Bin Boxes Desk Monitor
ORB-SLAM2 [50] I 2.5 7.0 9.3 10.3
DPVO [1] I 3.6 6.7 7.0 11.5
DPVO [1] E 5.8 7.2 7.8 14.4
EDS [2] I+E 2.6 5.8 5.0 8.0
DPVO [1] I+E 4.2 5.3 4.9 6.2
RAMP-VO (Ours) I+E 3.1 5.1 3.1 4.2
Results: In Table I, we start by analyzing the performance on the challenging Malapert landing dataset. RAMP-VO is able to recover accurate poses that deviate only $0.2\%$ to $1.7\%$ from the ground truth poses, if we consider that trajectories cover $250$km in distance. DPVO, instead, can not recover a valid trajectory leading to ATE errors of several kilometers, equivalent to $20\%$ to $30\%$. When events are added, DPVO decreases the error by $45.14\%$ to $29.46\%$, highlighting the importance of events in dark regions. We report a qualitative comparison on a Malapert sample in Figure 5.
On the Apollo landing dataset, both RAMP-VO versions, multi- and single-scale (SS), outperform image and image+event DPVO baselines by up to $30.77\%$, reaching an error from $2\%$ to $6\%$ of the ground truth. Contrary to Malapert, Apollo scenes feature clear frames and events, which explains the lower performance improvement given by the events. Single-scale and multi-scale RAMP-VO achieve similar results on Apollo, differently from Malapert where errors are around two times higher for single-scale. This suggests that when information is scarce, like in Malapert low-light environments, having the ability to focus on both global and fine-grained details improves robustness.
IV-C Comparison with State of the Art
TABLE III: Average Absolute Trajectory Error (m) on ECCV 2020 SLAM competition monocular test-split. Methods marked with (✓) use global optimization / loop closure. Top performing (non-global) method in bold, second best underlined.
@@ IV-C Comparison with State of the Art
TABLE IV: Avg. Absolute Trajectory Error (cm) and $R_{rmse}$ (deg) on EDS. ($*{}^{*}$) indicates methods trained on TartanAir train+test data. Top performing (non-global) VO method in bold, second best underlined.
@@
Results: Results for ECCV 2020 competition are available in Table III, where we follow the evaluation in [1], reporting the ATE[m] of the median of 5 runs with scale alignment. RAMP-VO is able to recover a better pose compared to all other image-based state-of-art methods [51, 54, 52, 3, 1] in most cases, outperforming by $20\%$ and $50\%$ DPVO [1] and DROID-SLAM using loop closure [3], respectively.
Results on Stereo DAVIS are reported in Table II. Given our emphasis on space applications, where space-graded cameras such as the AURICAMTM [55] often operate at low-FPS, we evaluate performance using the 5 FPS benchmark introduced in the supplementary analysis of [2]. Except for Bin, the proposed RAMP-VO outperforms all other baselines that use either both or just one of the two modalities on Stereo DAVIS, surpassing the top-performing method, EDS, by $25\%$. On the EDS benchmark reported in Table IV, our method consistently improves over DPVO and achieves an average 30% improvement over DEVO which, contrary to our method, is trained on additional data from the TartanAir test set. These benchmarks represent completely new scenarios compared to the TartanAir training setting. Stereo DAVIS features gray-scale frames and a lower $180\times 240$ resolution, while EDS has a higher resolution from Prophesee Gen3.1 and FLIR cameras, and an increased variety of test cases, including light and dark scenes and wider motions. Similar to the Apollo landing dataset, RAMP-VO is thus required to generalize from simulated to real sensor data.
@@ V Conclusion
In this work, we introduce RAMP-VO, an end-to-end VO system tailored for challenging environments such as those encountered during lunar descents. RAMP-VO exploits asynchronous and massively parallel encoders to efficiently and effectively fuse asynchronous event data into synchronous frames, achieving a 8×8$\times$8 × speedup and 33% improvement over state-of-the-art asynchronous encoders. Moreover, by incorporating events, RAMP-VO reduces the trajectory error of existing deep-learning-based solutions by 60%, as well as model-based VO methods fusing images and events by 27.6% on average. Experiments show that RAMP-VO can transfer zero-shot to real data despite being trained only on synthetic one, while still outperforming the other baselines. We designed RAMP-VO with a focus on space applications. While our primary focus was on improving accuracy and processing speed, future endeavors should prioritize optimizing the method for deployment on resource-constrained hardware. Approaches such as model quantization, network compression, and efficient hardware implementation should be explored to make our neural network space-ready [56]. Despite these limitations, we view this work as a milestone in event data fusion for VO, and we believe it can spark new interest in the use of event cameras and learning-based approaches for robust navigation.

### v3

@@ IV Experiments
Training. We train RAMP-VO on the synthetic TartanAir [5] dataset, with same train and test-split in [1], which we augment with events using VID2E [48]. During training, we follow the DPVO [1] training scheme to select and filter frames for training. Additionally, we also enforce a minimum of $1.2$M events between every pair of frames and remove additional frames if this condition is not satisfied. Finally, we feed RAMP-VO by interleaving two event stacks for every pair of selected frames. We create the first event stack by stacking the $600,000$ events received just before the mid-timestamp between the pair of images, and the second event stack by aggregating the $600,000$ events preceding the second frame. Since ground truth poses are only given for frames, we do not compute a loss for the mid-frame events. At inference time, we feed the model with asynchronous images and events, by creating a new event stack each time $M$ events are received. As the number of triggered events depends on the scene dynamics, with more events triggered with faster motion, the number of event stacks collected varies adaptively. For comparisons, we also test our model using the data feeding strategy used for training, which we call synchronized as events are collected at regular rates.
@@
Datasets.  We use the Stereo DAVIS [49] and EDS [2] datasets, and the ECCV 2020 SLAM competition TartanAir [5] test-split to benchmark our method. Stereo DAVIS and EDS are real-camera datasets, allowing us to test RAMP-VO’s robustness to domain shift and noise originating from the sensor or potential miscalibration. To ablate the crucial components of RAMP-VO architecture, we use instead the TartanAir test-split used in [1]. Additionally, we introduce two new benchmarks that replicate lunar surface landings and feature challenging lighting conditions: Apollo landing and Malapert landing. These benchmarks feature rapid motion, high dynamic range, and textureless terrains.
Malapert landing.  It features a total of $20$ minutes of simulated recordings from the Malapert south Moon region, divided into $2$ sequences. We use the planets and satellites simulator PANGU  [^3] https://pangu.software/ to generate realistic descent trajectories, each $250$ km long in altitude and $40$ km in translation, featuring partial or complete darkness. Ground truth poses of the spacecraft’s center of mass are provided at $5$Hz, together with $640\times 480$ synchronized RGB images. We generate synthetic events using Vid2E [48] with default settings.
Apollo landing.  A real dataset consisting of a total of $5$ minutes of recording, split into $6$ trajectories, featuring both vertical and lateral descent trajectories on a $260\times 260$ cm scale replica of the Apollo 17 landing site. Frames and events are recorded with a beam-splitter similar to the one used in [2]. Poses are recorded with an OptiTrack motion capture system. We downsample frames and events to a resolution of $640\times 480$ before processing.
@@ IV-A Effects of RAMP blocks
Results: From Figure 4 (a), DPVO+events has the lowest performance among the methods we tested (AUC of $0.56$), highlighting that a trivial event-image fusion is not sufficient to exploit events. Moreover, our RAMP-VO encoder in synchronous mode outperforms the baseline that uses RAM-Net both when we use multiple scales, as in RAM-Net, but also when just one scale is used. The single-scale RAMP-VO (AUC of $0.71$) achieves a 11% increase over the RAM-Net encoder (AUC $0.64$), while the multi-scale version (AUC $0.81$) improves by 27% over RAM-Net. When events are processed asynchronously, using $M=250000$, the performance increases even more, reaching a $0.85$ AUC and a 33% improvement.
@@
Low Framerate VO: In this section, we demonstrate that our method is not only able of asynchronous processing, but it can also cope with frames provided at very low rates. Since our RAMP encoders build sensor-agnostic features, we can exploit the faster data stream, i.e., event voxels, to keep the feature embedding updated, and generate a consistent pose.
To test this capability, we design an experiment on the carwelding hard p003 and carwelding easy p007 sequences of the TartanAir, where we artificially reduce the framerate of the images, subsampling them, but keeping event stacks fixed at 20 Hz. We then evaluate the trajectory error against ground truth poses at 20 Hz for DPVO+events, RAMP-VO, and RAMP-VO with RAM-Net-like encoder, and report the results in Fig. 4 (c). Our RAMP encoder consistently outperforms the baselines, including the RAM-Net-like encoder, indicating its superior performance on asynchronous data.
@@ IV-B Results on Space Data
DPVO [1] I 73.1 48.2 0.9 0.3 1.3
DPVO [1] I+E 40.1 34.0 0.9 0.2 1.1
RAMP-VO SS (Ours) I+E 1.1 9.4 0.7 0.2 1.0
RAMP-VO (Ours) I+E 0.6 4.3 0.8 0.2 0.9
TABLE II: Average Absolute Trajectory Error (m) on ECCV 2020 SLAM competition monocular test-split. Methods marked with (✓) use global optimization / loop closure. Top performing (non-global) method in bold, second best underlined.
@@
TABLE III: Average Absolute Trajectory Error (cm) on Stereo Davis
Input Bin Boxes Desk Monitor
ORB-SLAM2 [50] I 2.5 7.0 9.3 10.3
DPVO [1] I 3.6 6.7 7.0 11.5
DPVO [1] E 5.8 7.2 7.8 14.4
EDS [2] I+E 2.6 5.8 5.0 8.0
DPVO [1] I+E 4.2 5.3 4.9 6.2
RAMP-VO (Ours) I+E 3.1 5.1 3.1 4.2
Results: In Table I, we report performance on the Malapert landing dataset, where RAMP-VO recovers accurate poses deviating only 0.2% to 1.7% from the ground truth over 250km trajectories. DPVO, instead, can not recover a valid trajectory leading to ATE errors of several kilometers, equivalent to $20\%$ to $30\%$. When events are added, DPVO decreases the error by $45.14\%$ to $29.46\%$, highlighting the importance of events in dark regions. We report a qualitative comparison on a Malapert sample in Figure 5.
On the Apollo landing dataset, both RAMP-VO versions, multi- and single-scale (SS), outperform image and image+event DPVO baselines by up to $30.77\%$, reaching an error from $2\%$ to $6\%$ of the ground truth. Single-scale and multi-scale RAMP-VO achieve similar results on Apollo, contrary to Malapert’s low-light environments, where single-scale errors are twice as high, indicating the need to focus on both global and fine-grained details for improved robustness.
IV-C Comparison with State of the Art
TABLE IV: Avg. Absolute Trajectory Error (cm) and $R_{rmse}$ (deg) on EDS. () indicates methods trained on TartanAir train+test data. Top performing (non-global) VO method in bold, second best underlined.
@@ IV-C Comparison with State of the Art
Results: Results for ECCV 2020 competition are available in Table II, where we follow the evaluation in [1], reporting the ATE[m] of the median of 5 runs with scale alignment. RAMP-VO is able to recover a better pose compared to all other image-based state-of-art methods [51, 54, 52, 3, 1] in most cases, outperforming by $19\%$ and $48\%$ DPVO [1] and DROID-SLAM using loop closure [3], respectively.
Results on Stereo DAVIS are reported in Table III. Given our emphasis on space applications, where space-graded cameras such as the AURICAMTM  [^4] AURICAM’s datasheet available on www.sodern.com. often operate at low-FPS, we evaluate performance using the 5 FPS benchmark introduced in the supplementary analysis of [2]. Except for Bin, the proposed RAMP-VO outperforms all other baselines that use either both or just one of the two modalities on Stereo DAVIS, surpassing the top-performing method, EDS, by $27.6\%$. On the EDS benchmark reported in Table IV, our method consistently improves over DPVO and achieves an average 30% improvement over DEVO which, contrary to our method, is trained on additional data from the TartanAir test set. These benchmarks represent completely new scenarios compared to the TartanAir training setting. Stereo DAVIS features gray-scale frames and a lower $180\times 240$ resolution, while EDS has a higher resolution from Prophesee Gen3.1 and FLIR cameras, and an increased variety of test cases, including light and dark scenes and wider motions. Similar to the Apollo landing dataset, RAMP-VO is thus required to generalize from simulated to real sensor data.
@@ V Conclusion
In this work, we introduce RAMP-VO, an end-to-end VO system tailored for challenging environments such as those encountered during lunar descents. RAMP-VO uses RAMP encoders to fuse event data into frames, achieving 8×8$\times$8 × speedup and 33% improvement over state-of-the-art asynchronous encoders. Moreover, by incorporating events, RAMP-VO reduces the trajectory error of existing image-only deep-learning-based solutions by 58.8%, as well as event-only methods by 30.6% on average. Experiments show that RAMP-VO can transfer zero-shot to real data despite being trained only on synthetic one, while still outperforming the other baselines. We designed RAMP-VO for space applications focusing on improving accuracy and sensor fusion. Future efforts should prioritize optimizing the method for deployment on resource-constrained, space-graded hardware leveraging techniques such as model quantization, network compression, and efficient hardware implementations [55]. Despite these limitations, we view this work as a milestone in event data fusion for VO, and we believe it can spark new interest in the use of event cameras and learning-based approaches for robust navigation.